in the laboratory a student combines 27.2 ml of a .338 M barium bromide solution with 17.7 ml of a .566 M barium acetate solution, what is the final concentration of barium cation?


1
Expert's answer
2021-02-04T07:02:53-0500

Q160904


In the laboratory, a student combines 27.2 ml of a .338 M barium bromide solution with 17.7 ml of a .566 M barium acetate solution, what is the final concentration of barium cation?


Solution:


Barium bromide and barium acetate both are soluble in an aqueous medium.

We are given the concentration and volume of BaBr2 and Ba(C2H3O2)2 solution.

Using this we can find the moles of BaBr2 and Ba(C2H3O2)2.

From moles of BaBr2 and Ba(C2H3O2)2 we can find the moles of Ba+2 present in both the

compounds.

Total moles of Ba+2 after mixing will be equal to the sum of moles of Ba+2 present in both compounds.


Once we got the total moles of Ba+2, we can use the total moles of Ba+2 and total volume in the molarity formula and find the final concentration of Barium cation.


Step 1: To find moles of BaBr2


We are given 27.2 ml of a .338 M barium bromide solution.


27.2 mL in 'L' = "27.2\space mL * \frac{1\space L}{1000\space mL} = 0.0272\space L ;"



"molarity = \frac{ moles \space of\space BaBr_2}{volume \space of \space BaBr_2 \space in\space 'L' }"



"0.338\space M = \frac{ moles \space of\space BaBr_2}{0.0272\space L}"


multiply both the side by 0.0272L we have


"0.338\space mol/L * 0.0272\space L = \frac{ moles \space of\space BaBr_2}{0.0272\space L} * 0.0272\space L" ; ( since M = mol/L)


0.009194 mol = moles of BaBr2



Step 2: To find the moles of Ba(C2H3O2)2.


We are given 17.7 ml of a .566 M barium acetate solution.


17.7 mL in 'L' = "17.7\space mL * \frac{1\space L}{1000\space mL} = 0.0177\space L ;"



"molarity = \frac{ moles \space of\space Ba(C_2H_3O_2)_2}{volume \space of \space Ba(C_2H_3O_2)_2 \space in\space 'L' }"



"0.566\space M = \frac{ moles \space of\space Ba(C_2H_3O_2)_2}{0.0177\space L}"


multiply both the side by 0.0177L we have


"0.566\space mol/L * 0.0177\space L = \frac{ moles \space of\space Ba(C_2H_3O_2)_2}{0.0177\space L} * 0.0177\space L" ; ( since M = mol/L)


0.01002 mol = moles of Ba(C2H3O2)2.




Step 3: To find the total moles of Ba+2 ion.


The dissociation of BaBr2 in the solution is given as


BaBr2 ===> Ba+2 (aq) + 2 Br- (aq) ;


The mole to mole ratio of BaBr2 and Ba+2 in the reaction is 1 :1.

so moles of Ba+2 obtained from BaBr2 will be the same, that is 0.009194 mol Ba+2



The dissociation of Ba(C2H3O2)2. is given as


Ba(C2H3O2)2. ===> Ba+2 (aq) + 2C2H3O2-1 (aq)


The mole to mole ratio of Ba(C2H3O2)2. and Ba+2 in the reaction is 1 :1.

so moles of Ba+2 obtained from Ba(C2H3O2)2. will be the same, that is 0.01002 mol Ba+2




Total moles of Ba+2 = 0.009194 mol Ba+2 + 0.01002 mol Ba+2

= 0.019214 mol Ba+2



Step 4: To find the final concentration of Ba+2 after mixing.


Total volume of the solution = 27.2mL + 17.7 mL = 44.9 mL.


Final volume in 'L' = "44.9\space mL * \frac{1\space L}{1000\space mL} = 0.0449\space L ;"



Plug 0.019214 mol Ba+2 and 0.0449 L in the molarity formula and find the final concentration of

Ba+2 cation.


"molarity = \frac{ moles \space of\space Ba^{+2}}{volume \space of \space final \space solution \space in\space 'L' }"



"molarity = \frac{ 0.019214 \space mol \space of\space Ba^{+2}}{0.0449L} = 0.428mol/L ;"




Hence the final concentration of Ba+2 cation in the solution is 0.428 M or 0.428 mol/L.